# New Taipei City, 251301, Taiwan

### v1

@@ Abstract
Children often suffer wrist trauma in daily life, while they usually need radiologists to analyze and interpret X-ray images before surgical treatment by surgeons. The development of deep learning has enabled neural networks to serve as computer-assisted diagnosis (CAD) tools to help doctors and experts in medical image diagnostics. Since the You Only Look Once Version-8 (YOLOv8) model has obtained the satisfactory success in object detection tasks, it has been applied to various fracture detection. This work introduces four variants of Feature Contexts Excitation-YOLOv8 (FCE-YOLOv8) model, each incorporating a different FCE module (i.e., modules of Squeeze-and-Excitation (SE), Global Context (GC), Gather-Excite (GE), and Gaussian Context Transformer (GCT)) to enhance the model performance. Experimental results on GRAZPEDWRI-DX dataset demonstrate that our proposed YOLOv8+GC-M3 model improves the mAP@50 value from 65.78% to 66.32%, outperforming the state-of-the-art (SOTA) model while reducing inference time. Furthermore, our proposed YOLOv8+SE-M3 model achieves the highest mAP@50 value of 67.07%, exceeding the SOTA performance. The implementation of this work is available at https://github.com/RuiyangJu/FCE-YOLOv8.
@@
Deep Learning Computer Vision Object Detection Fracture Detection Medical Image Processing Medical Image Diagnostics You Only Look Once (YOLO) Feature Contexts Excitation
@@ 1 Intorduction
Wrist trauma is common among children [14, 35]. Meena et al. [32] have shown that most fractures occur around 2 centimeters from the distal radius near the joint. If they are not treated promptly and effectively, such trauma will result in wrist joint deformities, limited joint motion, and chronic pain [1]. In severe cases, an incorrect diagnosis would lead to lifelong complications and inconvenience [24].
@@
AI models play a crucial role in the medical field [30]. With the development of deep learning, neural networks are increasingly used as computer-assisted diagnosis (CAD) tools [25, 31, 39, 40] to aid doctors and experts in analyzing medical images (e.g., X-ray images) [47]. Object detection networks can accurately predict fractures and reduce the probability of misdiagnosis. Although two-stage object detection networks [9, 10, 37] obtain the excellent performance, one-stage object detection networks [36] offer faster inference time. With the continued release of You Only Look Once (YOLO) series models, YOLOv8 [19] and YOLOv9 [43] models can perform object detection efficiently on low-computing platforms [22].
@@
Since the release of the GRAZPEDWRI-DX [33] dataset by the Medical University of Graz, Ju et al. [21] have been the first to deploy YOLOv8 models to predict pediatric wrist fractures. This deployment helped surgeons to interpret fractures in X-ray images, reducing misdiagnosis and providing a better information base for the surgery. Chien et al. [5] developed YOLOv8-AM models by incorporating different attention modules [29, 44, 46, 48] into YOLOv8 [19], which improved the model performance in fracture detection. In addition, Chien et al. [4] applied YOLOv9 to this dataset, achieving the state-of-the-art (SOTA) model performance. However, there remains a great possibility for improving the performance of current SOTA models.
Models Params FLOPs mAPval 50
Inference
Time
YOLOv8+ResCBAM(SOTA) 53.87M 196.2G 65.78% 8.7ms
YOLOv8+GC-M3(Ours) 43.85M 165.6G 66.32% 7.9ms
YOLOv8+GCT-M2(Ours) 43.64M 165.4G 65.67% 15.3ms
YOLOv8+SE-M3(Ours) 43.85M 165.6G 67.07% 15.3ms
YOLOv8+GE-M2(Ours) 43.65M 165.4G 65.99% 15.3ms
@@
Figure 1: GRAZPEDWRI-DX Dataset mAP@50 vs. Inference Time. The input image size is 1024; the sizes of YOLOv8 and its improved models are all large (L), and the size of YOLOv9 model is extended (E). The inference time per image is measured on NVIDIA GeForce RTX 3090 GPUs. Our proposed YOLOv8+GE, YOLOv8+GC, and YOLOv8+SE models outperform the current SOTA model on mAP@50 value.
Studies [3, 16, 17, 38] have shown that incorporating different Feature Contexts Excitation (FCE) modules into Convolutional Neural Networks (CNNs) can enhance the model performance. For instance, the Squeeze-and-Excitation (SE) [17] module can reduce the Top-5 error from 7.48% to 6.62% for the ResNet-50 [13] model on the ImageNet dataset [6], with only a slight increase in computational cost. The Global Context (GC) [3] module can enhance ResNet-50 [13] model performance by 2.7% $\uparrow$ on $AP^{bbox}$ and 2.4% $\uparrow$ on $AP^{mask}$ for object detection on the COCO 2017 [27] validation set. The ResNet-50 [13] model with the Gather-Excite (GE) [16] module outperforms its 101-layer counterpart on the ImageNet dataset [6] without adding any learnable parameters. Incorporating the Gaussian Context Transformer (GCT) [38] module into Mask R-CNN [12] with ResNet-50 [13] model increases $AP^{bbox}_{0.5}$ from 38.2% to 39.4% on the COCO 2017 [27] validation set. Therefore, we propose FCE-YOLOv8 models for pediatric wrist fracture detection, each incorporating SE [17], GC [3], GE [16], and GCT [38] modules into YOLOv8, respectively, to enhance the model performance, and the results are shown in Fig. 1. This paper makes the following contributions:
  • (1)
    Introduces four different FCE modules to YOLOv8 model architecture, respectively, and proposes the FCE-YOLOv8 model, specifically designed for pediatric wrist fracture detection. One variant of our proposed FCE-YOLOv8 models, YOLOv8+SE, obtains mAP@50 values surpassing the SOTA level on the GRAZPEDWRI-DX public dataset.
  • (2)
    Since the main improvements in YOLOv8 can be categorized into three methods, we conduct the experiment to evaluate the most efficient method for each variant of our proposed FCE-YOLOv8 models that incorporates different FCE modules.
  • (3)
    Our proposed YOLO+GC model achieves the second-best performance on mAP@50 value on the GRAZPEDWRI-DX public dataset with only a slight increase in inference time, suitable as CAD tools to assist doctors and experts in interpreting pediatric wrist trauma X-ray images in real-world.
@@
The rest of this paper is organized as follows: Section 2 reviews various research on pediatric wrist fracture detection based on YOLO models and related modules. Section 3 introduces FCE-YOLOv8 model architectures on three different improved methods, and FCE modules used in this work. Section 4 presents the experiment conducted to evaluate the most efficient improved method for different FCE-YOLOv8 models, and compares our proposed models with other SOTA models. Section 5 discusses the limitations of this work, particularly the class imbalance in the used dataset. Finally, Section 6 summarises the outcomes of this paper and suggests directions for future work.
2 Related Work
2.1 YOLO Models for Pediatric Wrist Fracture Detection
2.2 Related Modules
FCE modules (i.e., SE [17], GC [3], GE [16], and GCT [38]) used in this work can enhance the feature modeling capabilities of neural networks in object detection tasks. (1) SE [17], introduced by Hu et al. in 2017, can improve the network ability to select channel features. The main contribution of this module is to dynamically adjust the channel weights of the feature map through the “Squeeze” and “Excitation” mechanisms. (2) GC [3], proposed by Cao et al. in 2019, can enhance the convolutional operations by efficiently capturing the global context information, which is the simplified and optimized version of Non-local (NL) [45] block. (3) GE [16], presented by Hu et al. in 2018, improves the SE module to better integrate the spatial and channel information. Compared to SE, GE captures more detailed spatial information and enables spatial adjustments at more granularity. (4) GCT [38], introduced by Ruan et al. in 2021, incorporates the Gaussian distribution into attention mechanism to capture the global context. This module uses the kernel function of Gaussian distribution to model the global context correlations, which further improves the accuracy of context capturing.
@@ 2.2 Related Modules
Figure 5: Detailed illustration of FCE modules used in this work: (a) SE module, (b) GC module, (c) GE module, and (d) GCT module.
@@ 3 Proposed Method
3.1 Network Architecture
The network architecture of YOLOv8 [19] comprises three components: Backbone, Neck, and Head. The Backbone component employs the Cross Stage Partial (CSP) [42] strategy, which divides the feature map into two parts. One part uses convolution operations, while the other is concatenated with the convolutional output of the first part. In the Neck component, YOLOv8 applies a combination of Feature Pyramid Network (FPN) [26] and Path Aggregation Network (PAN) [28] for multi-scale feature fusion. This component deletes convolution operations during the upsampling stage, allowing the top feature maps to retain more information by utilizing more layers, while minimizing the loss of location information in the bottom feature maps due to fewer convolutional layers. In addition, YOLOv8 employs a decoupled Head [8] to separate the classification and detection tasks. It replaces the anchor-based method used in YOLOv5 [20] with an anchor-free method [7], which locates objects by recognizing their center points and predicting the distance from the center to the bounding box.
To enhance global feature extraction and capture comprehensive information from medical images, we have designed three improved methods by incorporating different FCE modules into different components of YOLOv8 network architecture, as illustrated in Figs. 2, 3, and 4, respectively. Specifically, in the improved method 1 (M1), FCE modules (i.e., SE, GC, GE, and GCT) are added after the Spatial Pyramid Pooling - Fast (SPPF) [19] layer in the Backbone component. In the improved method 2 (M2), different FCE modules are integrated into the final C2f [19] module of the Head component. While in the improved method 3 (M3), FCE modules are added after each of the four C2f modules in the Head component. For each FCE-YOLOv8 model, we select the most appropriate method based on experimental results, as discussed in Section 4.5.
3.2 Feature Contexts Excitation Modules
This work uses four different FCE modules, including SE [17], GC [3], GE [16], and GCT [38].
@@ 3.2 Feature Contexts Excitation Modules
GE can efficiently aggregate the global context information while maintaining spatial resolution, without significantly increasing the model computational demands. This makes it possible for use in CAD tools on the low-computing-power platform for fracture prediction. By gathering the global context information from the entire feature map, GE captures broader context effectively. In addition, different from other methods [18, 41] that aggregate global context information, GE preserves spatial resolution, which is useful for tasks requiring precise localization, such as fracture detection. Overall, GE can enhance the feature representation capabilities of the network, improving YOLOv8 model performance.
GCT can enhance the ability of YOLOv8 to capture context information by incorporating the Gaussian filter. This integration improves the capacity of the model for data modeling and allows for more effective utilization of training data by weighting both local and global context information. In addition, this module increases the efficiency of data utilization, reducing the demands for extensive datasets. Furthermore, by replacing the traditional self-attention mechanism with the Gaussian filter, GCT decreases computational complexity and enhances efficiency, particularly with large-scale data. In general, the Gaussian filter contributes to the model stability during training, mitigating issues such as gradient explosion or vanishing gradients. This results in improved convergence speed and overall performance of YOLOv8.
Table 1: Experimental results of selecting improved methods for pediatric wrist fracture detection using the GRAZPEDWRI-DX dataset for YOLOv8+SE model. The best performance and the 2nd best performance are in red and blue colors, respectively.
Model Method
Model
Size
Input
Size
@@
Inference
Time
YOLOv8+SE M1 Small 640 11.30M 29.1G 62.18% 2.6ms
Medium 640 26.12M 79.8G 62.35% 5.1ms
Large 640 43.93M 166.3G 62.93% 7.6ms
YOLOv8+SE M2 Small 640 11.17M 28.7G 61.45% 2.5ms
Medium 640 25.90M 79.1G 61.75% 5.0ms
Large 640 43.67M 165.5G 64.08% 8.0ms
YOLOv8+SE M3 Small 640 11.19M 28.7G 61.72% 2.6ms
Medium 640 25.95M 79.2G 62.63% 5.0ms
Large 640 43.85M 165.6G 62.79% 7.6ms
YOLOv8+SE M1 Small 1024 11.30M 29.1G 63.24% 4.9ms
Medium 1024 26.12M 79.8G 63.70% 10.1ms
Large 1024 43.93M 166.3G 64.55% 15.4ms
YOLOv8+SE M2 Small 1024 11.17M 28.7G 62.90% 5.0ms
Medium 1024 25.90M 79.1G 63.68% 10.0ms
Large 1024 43.67M 165.5G 64.40% 14.7ms
YOLOv8+SE M3 Small 1024 11.19M 28.7G 62.77% 5.1ms
Medium 1024 25.95M 79.2G 64.10% 10.0ms
Large 1024 43.85M 165.6G 67.07% 15.3ms
Table 2: Experimental results of selecting improved methods for pediatric wrist fracture detection using the GRAZPEDWRI-DX dataset for YOLOv8+GC model. The best performance and the 2nd best performance are in red and blue colors, respectively.
Model Method
Model
Size
Input
Size
@@
Inference
Time
YOLOv8+GC M1 Small 640 11.34M 29.1G 60.68% 1.7ms
Medium 640 26.17M 79.9G 63.07% 2.5ms
Large 640 43.97M 166.3G 64.55% 3.5ms
YOLOv8+GC M2 Small 640 11.21M 28.7G 61.30% 1.9ms
Medium 640 25.95M 79.2G 62.71% 2.7ms
Large 640 43.70M 165.5G 64.54% 3.5ms
YOLOv8+GC M3 Small 640 11.24M 28.7G 62.19% 1.9ms
Medium 640 26.03M 79.2G 62.77% 2.7ms
Large 640 43.85M 165.6G 63.97% 3.6ms
YOLOv8+GC M1 Small 1024 11.34M 29.1G 63.24% 2.9ms
Medium 1024 26.17M 79.9G 63.87% 5.1ms
Large 1024 43.97M 166.3G 64.85% 7.8ms
YOLOv8+GC M2 Small 1024 11.21M 28.7G 63.51% 2.8ms
Medium 1024 25.95M 79.2G 65.25% 5.1ms
Large 1024 43.70M 165.5G 65.70% 7.8ms
YOLOv8+GC M3 Small 1024 11.24M 28.7G 64.16% 2.8ms
Medium 1024 26.03M 79.2G 64.33% 5.2ms
Large 1024 43.85M 165.6G 66.32% 7.9ms
Table 3: Experimental results of selecting improved methods for pediatric wrist fracture detection using the GRAZPEDWRI-DX dataset for YOLOv8+GE model. The best performance and the 2nd best performance are in red and blue colors, respectively.
Model Method
Model
Size
Input
Size

### v2

@@ Abstract
Children often suffer wrist trauma in daily life, while they usually need radiologists to analyze and interpret X-ray images before surgical treatment by surgeons. The development of deep learning has enabled neural networks to serve as computer-assisted diagnosis (CAD) tools to help doctors and experts in medical image diagnostics. Since YOLOv8 model has obtained the satisfactory success in object detection tasks, it has been applied to various fracture detection. This work introduces four variants of Feature Contexts Excitation-YOLOv8 (FCE-YOLOv8) model, each incorporating a different FCE module (i.e., modules of Squeeze-and-Excitation (SE), Global Context (GC), Gather-Excite (GE), and Gaussian Context Transformer (GCT)) to enhance the model performance. Experimental results on GRAZPEDWRI-DX dataset demonstrate that our proposed YOLOv8+GC-M3 model improves the mAP@50 value from 65.78% to 66.32%, outperforming the state-of-the-art (SOTA) model while reducing inference time. Furthermore, our proposed YOLOv8+SE-M3 model achieves the highest mAP@50 value of 67.07%, exceeding the SOTA performance. The implementation of this work is available at https://github.com/RuiyangJu/FCE-YOLOv8.
@@
deep learning, computer vision, object detection, fracture detection, medical image processing, medical image diagnostics, you only look once (yolo), feature contexts excitation
@@ 1 Intorduction
Wrist trauma is common among children [18]. If they are not treated promptly and effectively, such trauma will result in wrist joint deformities, limited joint motion, and chronic pain [1]. In severe cases, an incorrect diagnosis would lead to lifelong complications and inconvenience [15].
@@
With the development of deep learning, neural networks are increasingly used as computer-assisted diagnosis (CAD) tools to aid doctors and experts in analyzing medical images. Object detection networks can accurately predict fractures and reduce the probability of misdiagnosis. Although two-stage object detection networks obtain the excellent performance, one-stage object detection networks offer faster inference time. With the continued release of You Only Look Once (YOLO) series models, YOLOv8 [11] and YOLOv9 [21] models can perform object detection efficiently on low-computing platforms.
@@
Since the release of the GRAZPEDWRI-DX [16] dataset by the Medical University of Graz, Ju et al. [12] have been the first to deploy YOLOv8 models to predict pediatric wrist fractures. This deployment helped surgeons to interpret fractures in X-ray images, reducing misdiagnosis and providing a better information base for the surgery. Chien et al. [5] developed YOLOv8-AM models by incorporating different attention modules into YOLOv8 [11], which improved the model performance. In addition, Chien et al. [4] applied YOLOv9 to this dataset, achieving the state-of-the-art (SOTA) model performance. However, there remains a great possibility for improving the performance of current SOTA models.
This work proposes FCE-YOLOv8 models, incorporating SE [9], GC [3], GE [8], and GCT [19] modules into YOLOv8, respectively, to enhance the model performance, and the results are shown in Fig. 1. This paper makes the following contributions: 1) Introduces four different FCE modules to YOLOv8 model architecture, demonstrating significant improvements in model performance for pediatric wrist fracture detection. 2) Evaluates the most efficient model architecture method for each variant of the FCE-YOLOv8 models. 3) Our proposed YOLO+GC model achieves the second-best performance on mAP@50 value on the GRAZPEDWRI-DX public dataset with only a slight increase in inference time.
[…]
ResCBAM (SOTA) 65.78% 8.7ms
GC-M3 (Ours) 66.32% 7.9ms
GCT-M2 (Ours) 65.67% 15.3ms
SE-M3 (Ours) 67.07% 15.3ms
GE-M2 (Ours) 65.99% 15.3ms
@@
Figure 1: GRAZPEDWRI-DX Dataset mAP@50 vs. Inference Time. The input image size is 1024; the sizes of YOLOv8 and its improved models are all large (L), and the size of YOLOv9 model is extended (E).
@@
The rest of this paper is organized as follows: Section 2 reviews previous research on pediatric wrist fracture detection based on YOLO models. Section 3 introduces FCE-YOLOv8 model architectures on three different improved methods, and FCE modules used in this work. Section 4 presents the experiment conducted to evaluate the most efficient model architecture method for FCE-YOLOv8 models, and compares our proposed models with other SOTA models. Section 5 discusses the limitations of this work, particularly the class imbalance in the used dataset. Finally, Section 6 summaries the outcomes of this paper and suggests directions for future work.
@@ 2 Related Work
Figure 5: Diagram of proposed FCE modules: (a) SE module, (b) GC module, (c) GE module, and (d) GCT module.
@@ 3 Proposed Method
3.1 Baseline Model
The network architecture of YOLOv8 [11] comprises three components: Backbone, Neck, and Head. The Backbone component employs the Cross Stage Partial (CSP) [20] strategy, which divides the feature map into two parts. One part uses convolution operations, while the other is concatenated with the convolutional output of the first part. In the Neck component, YOLOv8 applies a combination of Feature Pyramid Network (FPN) and Path Aggregation Network (PAN) for multi-scale feature fusion. This component deletes convolution operations during the upsampling stage, allowing the top feature maps to retain more information by utilizing more layers, while minimizing the loss of location information in the bottom feature maps due to fewer convolutional layers.
To enhance global feature extraction and capture comprehensive information from medical images, we have designed three improved methods by incorporating different FCE modules into different components of YOLOv8 network architecture, as illustrated in Figs. 2, 3, and 4, respectively. Specifically, in the improved method 1 (M1), FCE modules (i.e., SE, GC, GE, and GCT) are added after the Spatial Pyramid Pooling - Fast (SPPF) [11] layer in the Backbone component. In the improved method 2 (M2), different FCE modules are integrated into the final C2f [11] module of the Head component. While in the improved method 3 (M3), FCE modules are added after each of the four C2f modules in the Head component.
3.2 FCE Modules
@@
GE can efficiently aggregate the global context information while maintaining spatial resolution, without significantly increasing the model computational demands. This makes it possible for use in CAD tools on the low-computing-power platform for fracture prediction. By gathering the global context information from the entire feature map, GE captures broader context effectively. In addition, different from other methods [10] that aggregate global context information, GE preserves spatial resolution, which is useful for tasks requiring precise localization, such as fracture detection. Overall, GE can enhance the feature representation capabilities of the network, improving YOLOv8 model performance.
GCT can enhance the ability of YOLOv8 to capture context information by incorporating the Gaussian filter. This integration improves the capacity of the model for data modeling and allows for more effective utilization of training data by weighting both local and global context information. In addition, this module increases the efficiency of data utilization, reducing the demands for extensive datasets. Furthermore, by replacing the traditional self-attention mechanism with the Gaussian filter, GCT decreases computational complexity and enhances efficiency, particularly with large-scale data. In general, the Gaussian filter contributes to the model stability during training, mitigating issues such as gradient explosion or vanishing gradients, which results in improved convergence model speed and performance.
4 Experiments
4.1 Datasets
4.2 Evaluation Metrics
Fracture detection on pediatric wrist X-ray images belongs to object detection tasks, and six evaluation metrics in object detection tasks are applicable, including the model parameters (Params), floating-point operations (FLOPs), F1-Score, mean average precision at 50% intersection over union (mAP@50), mean average precision from 50% to 95% intersection over union (mAP@50-95), and the inference time per image.
Params in the model are determined by the architecture layer count, the number of neurons per layer, and the overall complexity of the model. In general, an increase in Params corresponds to a larger model size. While larger models generally offer better model performance, they also require more computational resources. Therefore, it is important to balance the model size with computational cost in practical applications.
FLOPs representing floating-point operations, are commonly used to evaluate the computational complexity of the model. This metric serves as an indicator of the computational performance and speed of the models. In general, the models with lower FLOPs are more suitable for resource-limited environments, while the models with higher FLOPs require more powerful hardware support.
F1-Score is the harmonic mean of the Precision and Recall of the model, ranging from 0 to 1. The value closer to 1 indicates that the model has achieved a better balance between Precision and Recall. Conversely, if either Precision or Recall is close to 0, F1-Score will also approach 0, signaling poor model performance.
mAP is a widely used metric for evaluating the performance of object detection models. In object detection tasks, the model aims to recognize objects within an image and determines their locations. For each class, the model calculates the area under the Precision-Recall curve, known as average precision (AP), which measures the model detection performance for that specific class. The mAP score is calculated by averaging the AP values across all classes, providing an overall measure of the model detection accuracy.
Inference time denotes the time taken by the network model from the input of an X-ray image to the generation of the prediction output, encompassing preprocessing, inference, and post-processing steps. In this work, inference time per image was measured using a single NVIDIA GeForce RTX 3090 GPU.
Table 1: Experimental results of the SE module based on YOLOv8 on the GRAZPEDWRI-DX dataset. 1The input image size is 640. 2The input image size is 1024.
@@ 4.2 Evaluation Metrics
SE-M1-S 11.30M 29.1G 62.18% 2.6ms
SE-M1-M 26.12M 79.8G 62.35% 5.1ms
SE-M1-L 43.93M 166.3G 62.93% 7.6ms
SE-M2-S 11.17M 28.7G 61.45% 2.5ms
SE-M2-M 25.90M 79.1G 61.75% 5.0ms
SE-M2-L 43.67M 165.5G 64.08% 8.0ms
SE-M3-S 11.19M 28.7G 61.72% 2.6ms
SE-M3-M 25.95M 79.2G 62.63% 5.0ms
SE-M3-L 43.85M 165.6G 62.79% 7.6ms
SE-M1-S 11.30M 29.1G 63.24% 4.9ms
SE-M1-M 26.12M 79.8G 63.70% 10.1ms
SE-M1-L 43.93M 166.3G 64.55% 15.4ms
SE-M2-S 11.17M 28.7G 62.90% 5.0ms
SE-M2-M 25.90M 79.1G 63.68% 10.0ms
SE-M2-L 43.67M 165.5G 64.40% 14.7ms
SE-M3-S 11.19M 28.7G 62.77% 5.1ms
SE-M3-M 25.95M 79.2G 64.10% 10.0ms
SE-M3-L 43.85M 165.6G 67.07% 15.3ms
4.3 Data Preparation
To ensure fairness in model performance comparison, the same data is used for all model training. We randomly divide the GRAZPEDWRI-DX [16] dataset into training, validation, and test sets in the ratio of 70%, 20%, and 10%, respectively. Specifically, the training set contains 14,204 images (69.88%); the validation set contains 4,094 images (20.14%), and the test set contains 2,029 images (9.98%). Furthermore, data augmentation is performed on the training set before all models training. Specifically, this work adjusts the contrast and brightness of the X-ray images using the addWeighted function in the Open Source Computer Vision Library (OpenCV), and expanding the original training dataset from 14,204 images to 28,408 images.
Table 2: Experimental results of the GC module based on YOLOv8 on the GRAZPEDWRI-DX dataset. 1The input image size is 640. 2The input image size is 1024.
@@ 4.3 Data Preparation
GC-M1-S 11.34M 29.1G 60.68% 1.7ms
GC-M1-M 26.17M 79.9G 63.07% 2.5ms
GC-M1-L 43.97M 166.3G 64.55% 3.5ms
GC-M2-S 11.21M 28.7G 61.30% 1.9ms
GC-M2-M 25.95M 79.2G 62.71% 2.7ms
GC-M2-L 43.70M 165.5G 64.54% 3.5ms
GC-M3-S 11.24M 28.7G 62.19% 1.9ms
GC-M3-M 26.03M 79.2G 62.77% 2.7ms
GC-M3-L 43.85M 165.6G 63.97% 3.6ms
GC-M1-S 11.34M 29.1G 63.24% 2.9ms
GC-M1-M 26.17M 79.9G 63.87% 5.1ms
GC-M1-L 43.97M 166.3G 64.85% 7.8ms
GC-M2-S 11.21M 28.7G 63.51% 2.8ms
GC-M2-M 25.95M 79.2G 65.25% 5.1ms
GC-M2-L 43.70M 165.5G 65.70% 7.8ms
GC-M3-S 11.24M 28.7G 64.16% 2.8ms
GC-M3-M 26.03M 79.2G 64.33% 5.2ms
GC-M3-L 43.85M 165.6G 66.32% 7.9ms
4.4 Training
Table 3: Experimental results of the GE module based on YOLOv8 on the GRAZPEDWRI-DX dataset. 1The input image size is 640. 2The input image size is 1024.